OpenStudy (anonymous):

6b > or equal to 3 over 5

OpenStudy (anonymous):

\[6b \ge \frac{ 3 }{ 5 }\]

OpenStudy (zzr0ck3r):

6b>=3/5 multiply both sides by 1/6 b>= (1/6)*(3/5) = 1/10 b>= 1/10